OpenStudy (anonymous):

Verify the following results by direct integration laplace of ( e^-t*sin2t)= 2/s^2+2s+5

hartnn (hartnn):

where are u stuck ? u realize u need to use product rule for integration there ....

OpenStudy (anonymous):

yes im pretty much stuck ? product rule... i dont think you diffrentiate.

OpenStudy (anonymous):

you mean integration by parts?

hartnn (hartnn):

yes, i meant 'by parts'

OpenStudy (anonymous):

working on it ....

OpenStudy (anonymous):

i think you use the quotient rule not product rule .... any way i got -1/5*e^ -t*sin2t+2cos2t ,is that ron the right lines?

hartnn (hartnn):

where did s go ? didn't u realize u need to integrate \[\int\limits_{0}^{\infty}e^{-st}e^{-t}\sin2tdt\]

OpenStudy (anonymous):

can you use the laplace tables or do you have to do it from first principles?

hartnn (hartnn):

write that as \(e^{-t(s+1)}\) and then integrate by parts

hartnn (hartnn):

when u integrate by parts you should get \(e^{-t(s+1)}[-(s+1)sin2t-2cos2t]/(s^2+2s+5)\)

hartnn (hartnn):

tell me if u don't get that

OpenStudy (calculusfunctions):

First of all do you know Laplace transform? If not I can write up a very brief lesson to explain.

OpenStudy (anonymous):

Right i have the same ans as hartnn , but isn't the overall ans =2/s^2+2s+5...?

hartnn (hartnn):

yeah, now put the limits, first upper limit, t= infinity , then lower limit, t=0 u'll get the required answer

OpenStudy (anonymous):

right i got it ! finally thank you :)

hartnn (hartnn):

welcome ^_^